Please code in ARM cm4Problem 5 Write a program that corresponds to the following C code: if (a > b) c = c + 5; if (a < b) c = c - 5; if (a == b) c
Answer #1

      CMP A0, B0 ; compare a and b
      BEQ equal ; go to 'equal' if a=b
      BLT less ; go to 'less' if a<b
      ADD C0, C0, #5 ; otherwise: c=c+5
less  SUB C0, C0, #5 ; c=c-5
equal ADD C0, C0, #10 ; (else a must be equal to b, hence c=c+10)